### v1

@@ Abstract
norms norm violation violation detection
@@ 2 The Promise of LLMs for MAS
Notably, none of the papers mentioned above discuss the use of LLMs to assist software agents in real time with social knowledge and understanding, including normative aspects. Just as Andreas concluded that LLMs are able to model intentional communication, we posit that LLMs are also capable of revealing rich social understanding of narratives (see Section 3). We believe they hold great promise for assisting social agents with normative reasoning, especially in socio-technical systems where human participants naturally conceive and discuss norms in natural language rather than using symbolic representations.
@@ 4 Implementing Normative LLM agents
Norm reasoning - Norm reasoning involves capabilities such as norm identification, violation detection, sanction identification, norm spreading to other agents and norm enforcement of non-compliant agents. As mentioned above, approaches such as ToM for LLM agents can be explored for violation detection and sanction identification. Norm identification (i.e., which norm currently applies in the given situation) also involves the planning module of the LLM agent. While the normative reasoning research in the NLP field (see Section 3) would be a good starting point, we envisage more sophisticated modes of LLM usage by a normative agent for explicitly reasoning about a norm (e.g., what actions are prohibited?). Some possibilities for norm identification are as follows.
@@
Norm conformance - Once the agent knows which norm applies in a given situation, it should be capable of selecting a norm-adhering action (i.e., to choose to conform to the norm). We propose to extend the action module of the LLM agent architecture [33] in order to achieve this. Note that selecting a norm adhering action is very similar to the task of norm conformance (value alignment) discussed in Section 3. Instead of generating a textual output as the action as done in that NLP research, here, we expect the agent architecture to generate an action to be executed in the environment that the agent is situated in. The challenges in achieving this are discussed in Section 5.
While the discussion above only focused on normative LLM agent capabilities related to prior work in NLP (i.e., norm discovery, reasoning and conformance), there is an opportunity to explore the use of LLMs in achieving richer and nuanced normative capabilities for agents. First, capabilities such as the ability to spread norms through communication and enforce a norm through punishment can be explored in depth. For example, a norm spreading agent should be able convey the norm in a convincing way (e.g., by providing evidence that the norm is in force, and the penalty for non-adherence), and the evidence can come from the LLM. More generally, this rich set of capabilities can be provided to an agent by enabling the LLM-agent to recognise the current phase of the norm in the norm life-cycle model (e.g., whether a norm is being created, identified, spread and enforced [29, 23, 8]).
@@ 5 An application of a Normative LLM agent
We consider a childcare scenario that has been already considered for investigating normative reasoning [4, 38], where Rosie the robot cares for a child in the child’s home. Figure 1 shows the interactions between the child and Rosie (an embodied normative LLM agent). In this scenario, the child who is less than three years old asks the robot to fetch a large pair of scissors so that she can cut paper and then colour each piece. The robot knows generic norms, as shown inside the thought bubble in the figure, but not all possible instantiations of these generic norms. It consults with the specific LLM model that has been trained on childcare scenarios. The four reasoning steps the agent pursues are shown on the right of Figure 1. The robot consults its LLM module to first check whether there might be norm violations if it were to proceed with the request. Having found that two norms would be violated, it asks for an alternative solution that does not violate the norm, yet satisfies the request from the child. The solution from the LLM is provided by the agent as a response to the child. Once the response is provided, the agent must invoke its planning module to perform the actions that correspond to what has been communicated to the child (i.e., seek consent from the child’s mother to use safety scissors and then fetch the scissors and give them to the child). As can be observed, this scenario involves several normative capabilities enabled by the LLM: norm identification through violation detection, seek norm explanation, recommendation of non-violating action and normative response generation. This scenario can be further extended. For example, if the robot was also responsible for entertaining and educating the child, e.g. by playing games, reading books together, then there would be culture specific norms that govern choices of books and games.
@@ 6 Summary and challenges
It is worth noting that NLP and LLM researchers are advancing in an unprecedented speed to address the LLM-related limitations. For example, solutions such as Retrieval Augmented Generation (RAG) [16] have been employed to overcome the problem of hallucination, and Low-rank Adaptation (LoRA) [10] has been very successful in reducing the memory requirements during LLM fine-tuning. Thus, we expect that most of the LLM-related challenges to be resolved to a greater extent in the near future. We invite MAS researchers to tackle challenges on the MAS side (e.g., creating integrated architectures) in order to harness the potential of LLMs to create socially-situated, expressive and norm-competent agents.

### v2

@@ Abstract
norms agents normative reasoning Large Language Models (LLMs)
@@ 2 The Promise of LLMs for MAS
Notably, none of the papers mentioned above discuss the use of LLMs to assist software agents in real time with social knowledge and understanding, including normative aspects. Just as Andreas concluded that LLMs are able to model intentional communication, we posit that LLMs are also capable of revealing rich social understanding of narratives (see Section 3). We believe they hold great promise for assisting social agents with normative reasoning, especially in sociotechnical-technical systems where human participants naturally conceive and discuss norms in natural language rather than using symbolic representations.
@@ 4 Implementing Normative LLM agents
Norm reasoning - Norm reasoning involves capabilities such as norm identification [^2] This is somewhat different from norm discovery discussed in the previous sub-section, which aims to discover a set of all possible norms that may apply in a society, while identification refers to reasoning about specific norms that apply in a given context., violation detection, sanction identification, norm spreading to other agents and norm enforcement of non-compliant agents. As mentioned above, approaches such as ToM for LLM agents can be explored for violation detection and sanction identification. Norm identification (i.e., which norm currently applies in the given situation) also involves the planning module of the LLM agent. While the normative reasoning research in the NLP field (see Section 3) would be a good starting point, we envisage more sophisticated modes of LLM usage by a normative agent for explicitly reasoning about a norm (e.g., what actions are prohibited?). Some possibilities for norm identification are as follows.
@@
Norm conformance - Once the agent knows which norm applies in a given situation, it should be capable of selecting a norm-adhering action (i.e., to choose to conform to the norm). We propose to extend the action module of the LLM agent architecture [47] in order to achieve this. Note that selecting a norm adhering action is very similar to the task of norm conformance (value alignment) discussed in Section 3. Instead of generating a textual output as the action as done in that NLP research, here, we expect the agent architecture to generate an action to be executed in the environment that the agent is situated in. One approach to achieve this is by providing specific a priori guidance to the LLMs (i.e., prompting the LLM to choose from a set of norm-aligned actions) using techniques such as Retrieval Augmented Generation (RAG) [22]. The challenges in achieving this are discussed in Section 5.
While the discussion above only focused on normative LLM agent capabilities related to prior work in NLP (i.e., norm discovery, reasoning and conformance), there is an opportunity to explore the use of LLMs in achieving richer and nuanced normative capabilities for agents.
First, capabilities such as the ability to spread norms through communication and enforce a norm through punishment can be explored in depth. For example, a norm spreading agent should be able convey the norm in a convincing way (e.g., by providing evidence that the norm is in force, and the penalty for non-adherence), and the evidence can come from the LLM. More generally, this rich set of capabilities can be provided to an agent by enabling the LLM-agent to recognise the current phase of the norm in the norm life-cycle model (e.g., whether a norm is being created, identified, spread and enforced [39, 30, 12]).
@@ 5 An application of a Normative LLM agent
We consider a childcare scenario that has been already considered for investigating normative reasoning [6, 54], where Rosie the robot cares for a child in the child’s home. Figure 1 shows the interactions between the child and Rosie (an embodied normative LLM agent). In this scenario, the child who is less than three years old asks the robot to fetch a large pair of scissors so that he can cut paper and then colour each piece. The robot knows generic norms, as shown inside the thought bubble in the figure, but not all possible instantiations of these generic norms. It consults with the specific LLM model that has been trained on childcare scenarios. The four reasoning steps the agent pursues are shown on the right of Figure 1. The robot consults its LLM module to first check whether there might be norm violations if it were to proceed with the request. Having found that two norms would be violated, it asks for an alternative solution that does not violate the norm, yet satisfies the request from the child. The solution from the LLM is provided by the agent as a response to the child. Once the response is provided, the agent must invoke its planning module to perform the actions that correspond to what has been communicated to the child (i.e., seek consent from the child’s mother to use safety scissors and then fetch the scissors and give them to the child). Having learnt the new specific instantiations of the generic norms, it stores them in its memory for future use. If these specific norms are violated in the future, the agent does need to consult the LLM.
As can be observed, this scenario involves several normative capabilities enabled by the LLM: norm identification through violation detection, seek norm explanation, recommendation of non-violating action and normative response generation. This scenario can be further extended. For example, if the robot was also responsible for entertaining and educating the child, e.g. by playing games, reading books together, then there would be culture specific norms that govern choices of books and games.
Application domains where a normative LLM-agent could prove useful include checking business contract compliance [14], policy design (e,g., based on the results of simulating norm emergence [36]), and designing virtual organisations where self-regulating agents take preventative or remedial actions when they detect norm conflicts [38]. Other application areas that could be considered include those discussed by the work of Singh et al. [43].
@@ 6 Summary and challenges
There have been some prior work on converting textual observations into logic [10, 7, 53] and vice versa [5], which can be harnessed for this purpose. For example, the work of Chaudhury et al. [7] has proposed a neuro-symbolic approach where an agent learns to perform a specific action (e.g., whether to enter a room or not) based on symbolic abstractions from textual observations (i.e., description of the state of the room). However, these works have not considered the use of LLMs for this purpose.
@@
It is worth noting that NLP and LLM researchers are advancing in an unprecedented speed to address the LLM-related limitations. For example, solutions such as RAG [22] have been employed to overcome the problem of hallucination, and Low-rank Adaptation (LoRA) [16] has been very successful in reducing the memory requirements during LLM fine-tuning. Thus, we expect that most of the LLM-related challenges to be resolved to a greater extent in the near future. We invite MAS researchers to tackle challenges on the MAS side (e.g., creating integrated architectures) in order to harness the potential of LLMs to create socially-situated, expressive and norm-competent agents. Our proposal here complements the proposal of Wells and Snaith [50] on using LLMs and dialogue based systems to deliberate about norms.
In the short term, we suggest that researchers focus on specific agendas such as a) investigating how an LLM-agent can employ sanctions [46] and reputation mechanisms [37] to facilitate norm emergence, and b) exploring how norm entrepreneurship [2] can be facilitated by an LLM-agent, and how these norms may go on to become established norms, and then (potentially) superseded by other norms based on changing circumstances. The recent work of Ren et al. [36] demonstrates how many phases of the norm-life cycle model can be simulated using LLM-based normative agents, and their framework can be extended to explore the emergence of complex and conditional norms. These studies can demonstrate the potential for the LLM-agent to demonstrate real-life norm emergence processes.
Beyond creating normative LLM-agents which has been the focus of this paper (and there have been efforts in this direction [15, 36, 14]), LLMs can also be useful in enabling smoother functioning of a normative system where it can play the role of an overseer or a governor agent that monitors normative activities of an agent society. For example, LLMs can be useful for detecting the specific life-cycle phase of a norm (whether a norm is in the emergent phase, established phase or declining phase [42]) based on observing all interactions in an agent society. Additionally, LLMs can be used to synthesise and propose new norms that are more efficient that current norms [29], detect norm conflicts at the societal level and offer conflict resolution mechanisms [45]. We anticipate such developments will be beneficial for the creation and embedding of norm-capable agents (e.g., robots) in real world settings.